OpenStudy (anonymous):

solve: |4x-8|-3x=6x+2 and please write the solution set!!! thank you

OpenStudy (dumbcow):

get the absolute value part alone on one side --> |4x-8| = 9x+2 then split it into 2 cases, positive and negative --> 4x-8 = 9x+2 and --> 4x-8 = -(9x+2)

OpenStudy (anonymous):

and then please help me to solve it i tried but with the absolute value i got them wrong

OpenStudy (dumbcow):

well we got rid of the abs value can you solve the first case for x 4x-8 = 9x+2

OpenStudy (anonymous):

this is what i did: |4x-8|-3x=6x+2 4x-8=9x+2 -8=5x+2 -10=5x -10/5=5x/5 x=-2

OpenStudy (anonymous):

and when i plugged it in -2 works but the absolute value really kicks me off because when you solve it then it turns into the same thing if there werent any absolute value

OpenStudy (anonymous):

|4x-8|-3x=6x+2 |4(-2)-8|-3(-2)=6(-2)+2 |-8-8|-(-6)=-12+2 |-16|-(-6)=-10 |-16|+6=-10 |-10|=-10

OpenStudy (anonymous):

but how can they be the same when there is an absolute value to the left side?

OpenStudy (dumbcow):

that is correct...so this means that x=-2 is an "extraneous solution" , it doesn't work try the 2nd case

OpenStudy (anonymous):

4x-8=-(9x+2) 4x-8=-9x-2 13x-8=-2 13x=6 13x/6=6/6 x=2.166

OpenStudy (anonymous):

then that in turn creates a solution with a decimal so i dont think either works

OpenStudy (dumbcow):

almost...last step , divide by 13 13x = 6 x = 6/13 this should work when you plug it back in

OpenStudy (anonymous):

0.4615285?

OpenStudy (dumbcow):

yes, if you want the decimal form :)

OpenStudy (anonymous):

but will it work? because this is for alg 2.. and this is the only question that i am confused on

OpenStudy (dumbcow):

it will work |4(6/13) -8| = 9(6/13) +2 |24/13 - 8| = 54/13 + 2 |-80/13| = 80/13 80/13 = 80/13

OpenStudy (anonymous):

can you please explain how you did that?

OpenStudy (dumbcow):

i plugged in the solution x=6/13 into original equation |4x-8| = 9x+2 , to check that it would work for the fractions part --> 24/13 - 8 = 24/13 - 104/13 = -80/13 --> 54/13 +2 = 54/13 + 26/13 = 80/13

OpenStudy (anonymous):

so basically for the fractions you multiplied them?

OpenStudy (dumbcow):

yes...when multiplying fractions, top times top, bottom times bottom

OpenStudy (anonymous):

okay thank you

OpenStudy (dumbcow):

welcome